# Differential operators, anisotropy, and simplicial spheres

### v1

@@ Abstract.
We find identities involving differential operators in the generic artinian reduction of the Stanley–Reisner ring of a simplicial sphere in any positive characteristic. These identities generalize the characteristic $2$ identities used by Papadakis and Petrotou to give a proof of the algebraic $g$-conjecture. We use these identities to prove the anisotropy of certain forms on the generic artinian reduction of the Stanley–Reisner ring and to prove weak Lefschetz results.
@@ 1. Introduction
In [PP20], Papadakis and Petrotou found some identities involving differential operators in the generic artinian reduction of the Stanley–Reisner ring of a simplicial sphere over a field of characteristic $2$, and they used those identities to give a proof of the algebraic $g$-conjecture for simplicial spheres. They conjectured a generalization of their identities, and this conjecture was proven in [KX23]. These identities have enabled much recent progress in the study of simplicial spheres [APP21, LSN24, Oba24].
We prove analogous identities in the generic artinian reduction of the Stanley–Reisner ring of a simplicial sphere over a field of characteristic $p$ for any prime $p$. We use these identities to deduce algebraic properties of such generic artinian reductions, including proving conjectures of Adiprasito, Papadakis, and Petrotou. Our results hold for all connected oriented simplicial pseudomanifolds, and we work in that level of generality. Recall that a simplicial pseudomanifold of dimension $d-1$ is a pure simplicial complex $\Delta$ of dimension $d-1$ such that every $(d-2)$-dimensional face lies in exactly two facets, and each connected component of the geometric realization of $\Delta$ remains connected after we remove its $(d-3)$-skeleton.
@@
If $L=(\nu_{i,j})$ is an $m\times n$ matrix of nonnegative integers for some positive integer $m$, then define
@@
For example, if $F$ is a facet and $L$ is the $1\times n$ matrix whose $j$th entry is $1$ if $j\in F$ and 00 otherwise, then $x^{L}=x^{F}$. We use $(x^{L})^{1/p}$ to denote the monomial whose $p$th power is $x^{L}$ if such a monomial exists and zero otherwise. For each $i\in\{1,\dotsc,d\}$ and $j\in V$, the differential operator $\frac{\partial}{\partial a_{i,j}}$ acts on $K$. If $I=(I_{i,j})$ is a $d\times n$ matrix of nonnegative integers, define
@@
$\partial^{I}\deg(x^{J})=(-1)^{d}\deg((x^{I}\cdot x^{J})^{1/p})^{p},$
where $(x^{I}\cdot x^{J})^{1/p}$ is the $p$th root of $x^{I}\cdot x^{J}$ if $x^{I}\cdot x^{J}$ is a $p$th power, and is zero otherwise.
When $p=2$, the above theorem was conjectured in [PP20, Conjecture 14.1] and proved in [KX23, Theorem 4.1], and it generalizes [PP20, Proposition 5.1 and 5.7]. Our argument imitates the proof of [KX23, Theorem 4.1], but it requires some new ideas, see Remark 2.10.
@@
Suppose that $d=2$, $p=3$ and $F=\{1<2\}$ is a facet. Let $I=\begin{bmatrix}2&0&0&\cdots\\ 0&2&0&\cdots\\ \end{bmatrix}$ and $J=(1,1,0,\cdots)$. Then Theorem 1.1 says that $\partial^{I}\deg(x^{F})=\deg(x^{F})^{3}$. This corresponds to the fact that $\left(\frac{\partial}{\partial a_{1,1}}\right)^{2}\left(\frac{\partial}{\partial a_{2,2}}\right)^{2}\frac{1}{a_{1,1}a_{2,2}-a_{2,1}a_{1,2}}=\frac{1}{(a_{1,1}a_{2,2}-a_{2,1}a_{1,2})^{3}}$ over a field of characteristic $3$.
@@
In particular, Theorem 1.3 shows that if $g\in\overline{H}^{m}(\Delta)$ is nonzero for some $m\leq d/p$, then $g^{p}$ is nonzero. This proves [APP21, Conjecture 7.3]. Also, $\ell^{d-pm}\cdot g$ is nonzero, so the map from $\overline{H}^{m}(\Delta)$ to $\overline{H}^{d-(p-1)m}(\Delta)$ given by multiplication by $\ell^{d-pm}$ is injective. In particular, the map from $\overline{H}^{m}(\Delta)$ to $\overline{H}^{m+1}(\Delta)$ given by multiplication by $\ell$ is injective for $m\leq\frac{d-1}{p}$. This implies that the the weak Lefschetz property holds for $\overline{H}(\Delta)$ in degrees at most $\lfloor(d-1)/p\rfloor$, and $\dim\overline{H}^{0}(\Delta)\leq\dotsb\leq\dim\overline{H}^{\lfloor(d-1)/p\rfloor}(\Delta)$. Even for $\mathbb{Q}$-homology spheres, these inequalities cannot in general be proven using the characteristic $2$ results of [PP20] or [KX23], see Example 3.7. A proof of the strong Lefschetz property for $\overline{H}(\Delta)$ was announced in [APP21], see also [Adi18].
@@
Let $\Delta$ be a connected oriented simplicial $\mathbb{Z}$-homology manifold of dimension $d-1$ over a field $k$ of characteristic 00. Assume that the integral homology of $\Delta$ is torsion-free. For any nonzero $g\in\overline{H}^{m}(\Delta)$ with $m\leq d/t$, $\ell^{d-tm}\cdot g^{t}$ is nonzero.
@@
Our paper is organized as follows. In Section 2, we prove Theorem 1.1. In Section 3, we prove Theorem 1.3 and Theorem 1.4. See [KX23, Section 2] for background on pseudomanifolds, orientations, and degree maps.
@@ Acknowledgements
This work was conducted while the second and third authors were at the Institute for Advanced Study, where the third author received support from the Charles Simonyi endowment. The first author was supported by a NSERC Discovery grant.
2. Proof of the identity
@@ 2.1. Reductions
Theorem 1.1 for all $\Pi_{i}$ implies it for $\Delta$.
@@
We check that the corresponding summands on both sides are equal. If $x^{J}$ involves a variable $x_{j}$ such that $j$ is not a vertex of $\Pi_{i}$, then $\deg_{\Pi_{i}}$ vanishes on both sides. Similarly, if $x^{I}$ involves such a variable $x_{j}$, then the term on the right hand side vanishes; on the left hand side, $\deg_{\Pi_{i}}$ does not depend on variables $a_{i,j}$, so the derivative vanishes. In the case when $x^{I}x^{J}$ only involves variables $x_{j}$ such that $j$ is a vertex of $\Pi_{i}$, the two sides are equal by Theorem 1.1 applied to $\Pi_{i}$. ∎
@@
Let $I$ be a $d\times(d+1)$ matrix of nonnegative integers such that all row sums are $p-1$. Let $J$ be a $1\times(d+1)$ matrix of nonnegative integers whose row sum is $-1\pmod{p}$. Then
@@
In characteristic $p$, partial derivatives commute with $p$th powers. Multiplying both sides by $(X_{0}\cdots X_{d})^{p}$ gives the equivalent statement
$\partial^{I}X^{J}(X_{0}\cdots X_{d})^{p-1}=\begin{cases}(-1)^{d}X^{pL}&\textrm{ if }X^{I}X^{J}=X^{pL},\\ 0&\textrm{ otherwise.}\end{cases}$
The above equation is the statement of Theorem 2.3 for $I$ and $J^{\prime}$ such that $X^{J^{\prime}}=X^{J}(X_{0}\cdots X_{d})^{p-1}$. Note that the row sum of $J^{\prime}$ is $d+(p-1)(d+1)=pd+(p-1)$, which is $-1\pmod{p}$. ∎
@@ 2.2. $\operatorname{SL}(d,k)$-invariance
Let $B\in SL(d,k)$. Multiplication by $B$ from the left $A\mapsto B\cdot A$ gives a linear change of variables on the $a_{i,j}$, so it gives an action of $SL(d,k)$ on the polynomial ring $k[a_{i,j}]$. If $k$ is infinite, then the invariants of this action are the subring $k[X_{j}]\subseteq k[a_{i,j}]$. If $k$ is finite, then this subring is the set of absolute invariants, i.e., the polynomials in $k[a_{i,j}]$ that are invariant under the action of $SL(d,\overline{k})$, where $\overline{k}$ is the algebraic closure of $k$. The following lemma generalizes [KX23, Lemma 4.8].
@@
In the setup of Theorem 2.3, $\partial^{I}X^{J}\in k[a_{i,j}]$ is $\operatorname{SL}(d,k)$-invariant for any field $k$ of characteristic $p$. In particular, $\partial^{I}X^{J}\in\mathbb{F}_{p}[a_{i,j}]$ is an absolute $\operatorname{SL}(d,\mathbb{F}_{p})$-invariant.
@@
where $\partial^{I}(a^{\prime})$ is the derivative $\partial^{I}$ with each $\frac{\partial}{\partial a_{2,m}}$ replaced by $\frac{\partial}{\partial a_{2,m}}+c\frac{\partial}{\partial a_{1,m}}$. The derivative $\partial^{I}(a^{\prime})$ has one term equal to $\partial^{I}$. All other terms contain at least $p$ derivatives of the form $\frac{\partial}{\partial a_{1,m}}$. Lemma 2.6 below shows that all these other terms vanish, so
@@
where $b_{1}=(-1)^{j}Y_{2j}/Y_{12}$ and $b_{2}=(-1)^{j}Y_{1j}/Y_{12}$ are elements of $K$ that are independent of the variables $\{a_{r,j}:0\leq j\leq d\}$. Substituting the expression from (3) into $X^{J}$ for all $j>2$, we reduce to the case when $X^{J}=X_{1}^{m_{1}}X_{2}^{m_{2}}$ for some nonnegative integers $m_{1},m_{2}$ such that $m_{1}+m_{2}$ is the row sum of $J$. Since the row sum of $J$ is $-1\pmod{p}$ and is strictly greater than $p-1$, it must be at least $12p-$ 1. Hence either $m_{1}\geq p$ or $m_{2}\geq p$. Without loss of generality, assume that $m_{1}\geq p$. Then we can write $X^{J}=X_{1}^{p}X^{J^{\prime}}$ with the row sum of $J^{\prime}$ equal to $-1\pmod{p}$ and strictly less than the row sum of $J$. By induction, and since derivatives commute with $p$th powers, we have
@@
It follows that $\partial^{I}X^{J}$ is a constant multiple of the monomial $X_{M}$. We need to prove that this constant is equal to $(-1)^{d}$ if $(X^{M})^{1/p}$ exists, and it is zero otherwise.
@@
Lemma 2.6 gives that the partial derivative of $\partial^{I}X^{J}$ with respect to any $a_{i,j}$ is zero. This implies that $\partial^{I}X^{J}$ is a $p$th power in $k[a_{i,j}]$, and since the $X_{j}$ are irreducible, it is also a $p$th power in $k[X_{0},\ldots,X_{d}]$. ∎
@@
We have seen that $\partial^{I}X^{J}$ is a constant multiple of the monomial $X_{M}=X^{pL}$. We consider $I,J,L$ satisfying the equation
@@
Recall that $X_{j}$ is $(-1)^{j}$ times the determinant of the matrix $A$ with column $j$ removed. We write $\overline{X}_{j}$, $j=0,\ldots,d-1,$ for similar signed determinants in dimension $d-1$. Let $\overline{J}$ be the vector $J$ with its last entry removed. Let $I_{i}$ be the $i$th row of $I$, and let $\overline{I}_{i}$ be the row $I_{i}$ with its last entry removed. The vectors $\overline{I}_{i}$ for $i\neq r$ form the rows of $\overline{I}_{\hat{r}}$. Similarly, let $\overline{L}$ be $L$ with its last entry removed.

### v2

@@ Abstract.
We find identities involving differential operators in the generic artinian reduction of the Stanley–Reisner ring of a simplicial sphere in any positive characteristic. These identities generalize the characteristic $2$ identities used by Papadakis and Petrotou to give a proof of the algebraic $g$-conjecture. We show that these identities are a shadow of an identity on the degree map, and we use them to prove the anisotropy of certain forms on the generic artinian reduction of the Stanley–Reisner ring and to prove weak Lefschetz results.
@@ 1. Introduction
In [PP20], Papadakis and Petrotou found some identities involving differential operators in the generic artinian reduction of the Stanley–Reisner ring of a simplicial sphere over a field of characteristic $2$, and they used those identities to give a proof of the algebraic $g$-conjecture for simplicial spheres. They conjectured a generalization of their identities, and this conjecture was proven in [KX23]. These identities have enabled much recent progress in the study of simplicial spheres [APP21, LNS24, Oba24].
We show that these identities are a shadow of an identity on the degree map that we prove in Theorem 1.1 below. Furthermore, we generalize this identity to generic artinian reductions of Stanley–Reisner rings of simplicial spheres over a field of characteristic $p$ for any prime $p$. We use these identities to deduce algebraic properties of such generic artinian reductions, including proving conjectures of Adiprasito, Papadakis, and Petrotou. Our results hold for all connected oriented simplicial pseudomanifolds, and we work in that level of generality. Recall that a simplicial pseudomanifold of dimension $d-1$ is a pure simplicial complex $\Delta$ of dimension $d-1$ such that every $(d-2)$-dimensional face lies in exactly two facets, and each connected component of the geometric realization of $\Delta$ remains connected after we remove its $(d-3)$-skeleton.
@@
Let $L=(\nu_{i,j})$ be an $m\times n$ matrix of nonnegative integers for some positive integer $m$, i.e., the columns of $L$ are indexed by the vertices of $\Delta$. Set
@@
For example, if $F$ is a facet and $L$ is the $1\times n$ matrix whose $j$th entry is $1$ if $j\in F$ and 00 otherwise, then $x^{L}=x^{F}$. We use $(x^{L})^{1/p}$ to denote the monomial whose $p$th power is $x^{L}$ if such a monomial exists and zero otherwise. If $I=(I_{i,j})$ is a $d\times n$ matrix of non-negative integers, then set
$a^{I}\coloneqq\prod_{i,j}a_{i,j}^{I_{i,j}}\text{ and }I!\coloneqq\prod_{i,j}(I_{i,j})!.$
Theorem 1.1.
Let $\Delta$ be a connected oriented simplicial pseudomanifold of dimension $d-1$ over a field $k$ of characteristic $p$ with vertex set $V=\{1,\ldots,n\}$. Let $J$ be a $1\times n$ matrix of nonnegative integers with row sum $d$. Then
$\deg(x^{J})=(-1)^{d}\sum_{I}\deg((x^{I}\cdot x^{J})^{1/p})^{p}\frac{a^{I}}{I!},$
where $I$ varies over all $d\times n$ matrices of nonnegative integers whose row sums are all $p-1$.
In a different context, a similar identity appeared in [APPS22]. For each $i\in\{1,\dotsc,d\}$ and $j\in V$, the differential operator $\frac{\partial}{\partial a_{i,j}}$ acts on $K$. If $I=(I_{i,j})$ is a $d\times n$ matrix of nonnegative integers, define
@@
In characteristic $p$, partial derivatives commute with $p$th powers. If all row sums of $I$ are $p-1$, then after applying $\partial^{I}$ to the identity in Theorem 1.1, exactly one term on the right hand side survives, giving the following result.
@@
$\partial^{I}\deg(x^{J})=(-1)^{d}\deg((x^{I}\cdot x^{J})^{1/p})^{p}.$
In fact, we will deduce Theorem 1.1 from Theorem 1.2, showing that, in fact, Theorem 1.1 and Theorem 1.2 are equivalent. When $p=2$, the above theorem was conjectured in [PP20, Conjecture 14.1] and proved in [KX23, Theorem 4.1], and it generalizes [PP20, Proposition 5.1 and 5.7].
@@
Suppose that $d=2$, $p=3$, and $F=\{1<2\}$ is a facet. Let $I=\begin{bmatrix}2&0&0&\cdots\\ 0&2&0&\cdots\\ \end{bmatrix}$ and $J=(1,1,0,\cdots)$. Then Theorem 1.2 says that $\partial^{I}\deg(x^{F})=\deg(x^{F})^{3}$. This corresponds to the fact that $\left(\frac{\partial}{\partial a_{1,1}}\right)^{2}\left(\frac{\partial}{\partial a_{2,2}}\right)^{2}\frac{1}{a_{1,1}a_{2,2}-a_{2,1}a_{1,2}}=\frac{1}{(a_{1,1}a_{2,2}-a_{2,1}a_{1,2})^{3}}$ over a field of characteristic $3$.
@@
In particular, Theorem 1.4 shows that if $g\in\overline{H}^{m}(\Delta)$ is nonzero for some $m\leq d/p$, then $g^{p}$ is nonzero. This proves [APP21, Conjecture 7.3]. Also, $\ell^{d-pm}\cdot g$ is nonzero, so the map from $\overline{H}^{m}(\Delta)$ to $\overline{H}^{d-(p-1)m}(\Delta)$ given by multiplication by $\ell^{d-pm}$ is injective. In particular, the map from $\overline{H}^{m}(\Delta)$ to $\overline{H}^{m+1}(\Delta)$ given by multiplication by $\ell$ is injective for $m\leq\frac{d-1}{p}$. This implies that the weak Lefschetz property holds for $\overline{H}(\Delta)$ in degrees at most $\lfloor(d-1)/p\rfloor$, and $\dim\overline{H}^{0}(\Delta)\leq\dotsb\leq\dim\overline{H}^{\lfloor(d-1)/p\rfloor}(\Delta)$. Even for $\mathbb{Q}$-homology spheres, these inequalities cannot in general be proven using the characteristic $2$ results of [PP20] or [KX23], see Example 3.7. A proof of the strong Lefschetz property for $\overline{H}(\Delta)$ was announced in [APP21], see also [Adi18].
@@
Let $\Delta$ be a connected oriented simplicial $\mathbb{Z}$-homology manifold of dimension $d-1$ over a field $k$ of characteristic 00. Assume that the integral homology of $\Delta$ is torsion-free. For any $t\geq 2$ and nonzero $g\in\overline{H}^{m}(\Delta)$ with $m\leq d/t$, $\ell^{d-tm}\cdot g^{t}$ is nonzero.
@@
Our paper is organized as follows. In Section 2, we prove Theorem 1.2 and then use it to deduce Theorem 1.1. In Section 3, we prove Theorem 1.4 and Theorem 1.5. See [KX23, Section 2] for background on pseudomanifolds, orientations, and degree maps.
@@ Acknowledgements
This work was conducted while the second and third authors were at the Institute for Advanced Study, where the third author received support from the Charles Simonyi endowment. The first author was supported by a NSERC Discovery grant. We thank the referees for their helpful comments.
2. Proof of the identities
We begin by proving Theorem 1.2. Our argument imitates the proof of [KX23, Theorem 4.1], but it requires some new ideas, see Remark 2.10.
@@ 2.1. Reductions
If the conclusion of Theorem 1.2 holds for all $\Pi_{i}$, then it holds for $\Delta$.
@@
Using the conclusion of Theorem 1.2 for each $\Pi_{i}$, we check that the corresponding summands on both sides are equal. If $x^{J}$ involves a variable $x_{j}$ such that $j$ is not a vertex of $\Pi_{i}$, then $\deg_{\Pi_{i}}$ vanishes on both sides. Similarly, if $x^{I}$ involves such a variable $x_{j}$, then the term on the right hand side vanishes; on the left hand side, $\deg_{\Pi_{i}}$ does not depend on variables $a_{i,j}$, so the derivative vanishes. In the case when $x^{I}x^{J}$ only involves variables $x_{j}$ such that $j$ is a vertex of $\Pi_{i}$, the two sides are equal by Theorem 1.2 applied to $\Pi_{i}$. ∎
@@
Let $I$ be a $d\times(d+1)$ matrix of nonnegative integers such that all row sums are $p-1$. Let $J$ be a $1\times(d+1)$ matrix of nonnegative integers whose row sum is congruent to $-1\pmod{p}$. Then
@@
As partial derivatives commute with $p$th powers, multiplying both sides by $(X_{0}\cdots X_{d})^{p}$ gives the equivalent statement
$\partial^{I}\left(X^{J}(X_{0}\cdots X_{d})^{p-1}\right)=\begin{cases}(-1)^{d}X^{pL}&\textrm{ if }X^{I}X^{J}=X^{pL},\\ 0&\textrm{ otherwise.}\end{cases}$
The above equation is the statement of Theorem 2.3 for $I$ and $J^{\prime}$ such that $X^{J^{\prime}}=X^{J}(X_{0}\cdots X_{d})^{p-1}$. Note that the row sum of $J^{\prime}$ is $d+(p-1)(d+1)=pd+(p-1)$, which is congruent to $-1\pmod{p}$. ∎
@@ 2.2. $\operatorname{SL}(d,k)$-invariance
Let $B\in SL(d,k)$. Multiplication by $B$ from the left $A\mapsto B\cdot A$ gives a linear change of variables on the $a_{i,j}$, so it gives an action of $SL(d,k)$ on the polynomial ring $k[a_{i,j}]$. If $k$ is infinite, then the invariants of this action are the subring $k[X_{j}]$ ([Pro07, Theorem 13.5.5]). Theorem 2.3 depends only on the characteristic of the field $k$, not on the field itself. It suffices to prove it when $k$ is infinite.
The following lemma generalizes [KX23, Lemma 4.8].
@@
In the setup of Theorem 2.3, $\partial^{I}X^{J}\in k[a_{i,j}]$ is $\operatorname{SL}(d,k)$-invariant for any field $k$ of characteristic $p$.
@@
where $\partial^{I}(a^{\prime})$ is the derivative $\partial^{I}$ with each $\frac{\partial}{\partial a_{2,m}}$ replaced by $\frac{\partial}{\partial a_{2,m}}+c\frac{\partial}{\partial a_{1,m}}$, i.e.,
$\partial^{I}(a^{\prime})\coloneqq\prod_{m=0}^{d}\left(\frac{\partial}{\partial a_{2,m}}+c\frac{\partial}{\partial a_{1,m}}\right)^{I_{2,m}}\prod_{\begin{subarray}{c}i,m\\ i\not=2\end{subarray}}\left(\frac{\partial}{\partial a_{i,m}}\right)^{I_{i,m}}.$
The derivative $\partial^{I}(a^{\prime})$ has one term equal to $\partial^{I}$. All other terms contain at least $p$ derivatives of the form $\frac{\partial}{\partial a_{1,m}}$. Lemma 2.6 below shows that all these other terms vanish, so
@@
where $b_{1}=(-1)^{j}Y_{2j}/Y_{12}$ and $b_{2}=(-1)^{j}Y_{1j}/Y_{12}$ are elements of $K$ that are independent of the variables $\{a_{r,j}:0\leq j\leq d\}$. Substituting the expression from (3) into $X^{J}$ for all $j>2$, we reduce to the case when $X^{J}=X_{1}^{m_{1}}X_{2}^{m_{2}}$ for some nonnegative integers $m_{1},m_{2}$ such that $m_{1}+m_{2}$ is the row sum of $J$. Since the row sum of $J$ is $-1\pmod{p}$ and is strictly greater than $p-1$, it must be at least $12p-$ 1. Hence either $m_{1}\geq p$ or $m_{2}\geq p$. Without loss of generality, assume that $m_{1}\geq p$. Then we can write $X^{J}=X_{1}^{p}X^{J^{\prime}}$ with the row sum of $J^{\prime}$ congruent to $-1\pmod{p}$ and strictly less than the row sum of $J$. By induction, and since derivatives commute with $p$th powers, we have
@@
It follows that $\partial^{I}X^{J}$ is a constant multiple of the monomial $X^{M}$. We need to prove that this constant is equal to $(-1)^{d}$ if $(X^{M})^{1/p}$ exists, and it is zero otherwise.
@@
Lemma 2.6 gives that the partial derivative of $\partial^{I}X^{J}$ with respect to any $a_{i,j}$ is zero. This implies that $\partial^{I}X^{J}$ is a $p$th power in $k[a_{i,j}]$, and since the $X_{j}$ are irreducible, it is also a $p$th power in $k[X_{0},\ldots,X_{d}]$. As $\partial^{I}X^{J}$ is a constant multiple of $X^{M}$, we see that the constant must be 00 in order for it to be a $p$th power. ∎
@@
We have seen that $\partial^{I}X^{J}$ is a constant multiple of the monomial $X^{M}=X^{pL}$. We consider $I,J,L$ satisfying the equation
@@
Recall that $X_{j}$ is $(-1)^{j}$ times the determinant of the matrix $A$ with column $j$ removed. We write $\overline{X}_{j}$, $j=0,\ldots,d-1,$ for similar signed determinants in dimension $d-1$, i.e., $(-1)^{j}$ times the determinant of the result of removing the $j$th column, the last column, and the $r$th row from $A$. Let $\overline{J}$ be the vector $J$ with its last entry removed. Let $I_{i}$ be the $i$th row of $I$, and let $\overline{I}_{i}$ be the row $I_{i}$ with its last entry removed. The vectors $\overline{I}_{i}$ for $i\neq r$ form the rows of $\overline{I}_{\hat{r}}$. Similarly, let $\overline{L}$ be $L$ with its last entry removed.
@@
2.4. Proof of Theorem 1.1
We now deduce Theorem 1.1 from Theorem 1.2. For any $r\geq 1$, consider the differential operator
$\Psi_{r}\coloneqq\sum_{I}\frac{a^{I}}{I!}\partial^{I},$
where $I$ varies over all $d\times n$ matrices of nonnegative integers such that all row sums are equal to $r$. We prove an identity involving $\Psi_{r}$ which is valid in any characteristic.
Theorem 2.11.
Let $\Delta$ be a connected oriented simplicial pseudomanifold of dimension $d-1$ over a field $k$ with vertex set $V=\{1,\ldots,n\}$. Let $J$ be a $1\times n$ matrix of nonnegative integers with row sum $d$. Then
$\Psi_{r}\deg(x^{J})=(-1)^{dr}(r!)^{2d}\deg(x^{J}).$
Proof of Theorem 1.1.
Applying $\Psi_{p-1}$ to $\deg(x^{J})$, we have
$\begin{split}\Psi_{p-1}\deg(x^{J})&=\sum_{I}\frac{a^{I}}{I!}\partial^{I}\deg(x^{J})\\ &=(-1)^{d}\sum_{I}\deg((x^{I}\cdot x^{J})^{1/p})^{p}\frac{a^{I}}{I!},\end{split}$
where $I$ varies over all $d\times n$ matrices of nonnegative integers such that all row sums are equal to $p-1$, and the second equality is by Theorem 1.2. As $\Psi_{p-1}\deg(x^{J})=(-1)^{d(p-1)}((p-1)!)^{2d}\deg(x^{J})$ by Theorem 2.11 and $(-1)^{d(p-1)}((p-1)!)^{2d}\equiv 1\pmod{p}$, the result follows. ∎
In order to prove Theorem 2.11, we introduce some auxiliary differential operators that will allow us to factor $\Psi_{r}$. For $1\leq i\leq d$, define
$\Phi_{i,r}\coloneqq\sum_{1\leq j_{1},\dotsc,j_{r}\leq n}a_{i,j_{1}}\dotsb a_{i,j_{r}}\frac{\partial}{\partial a_{i,j_{1}}}\dotsb\frac{\partial}{\partial{a_{i,j_{r}}}}.$
Note that if $i\not=i^{\prime}$, then $\Phi_{i,r}$ and $\Phi_{i^{\prime},r}$ commute, as they involve distinct variables. We have that
(7) $\Psi_{r}=(r!)^{d}\prod_{i=1}^{d}\Phi_{i,r}.$
For $1\leq i\leq d$, define $\phi_{i}\coloneqq\sum_{j=1}^{n}a_{i,j}\frac{\partial}{\partial{a_{i,j}}}.$ Then we have the following identity.
Lemma 2.12.
For any $1\leq i\leq d$ and any $r\geq 1$, we have
$\Phi_{i,r}=\phi_{i}(\phi_{i}-1)\dotsb(\phi_{i}-(r-1)).$
Proof.
To simplify notation, let $\partial_{j}=a_{i,j}\frac{\partial}{\partial a_{i,j}}$. Note that the $\partial_{j}$ commute with each other. The one-variable case, i.e., that
$a_{i,j}^{\alpha}\left(\frac{\partial}{\partial a_{i,j}}\right)^{\alpha}=\partial_{j}(\partial_{j}-1)\cdots(\partial_{j}-(\alpha-1))$
for an integer $\alpha\geq 0$, can be checked by applying both sides to monomials $a_{i,j}^{\beta}$.
We will further use the notation
$[x]_{\alpha}=x(x-1)\cdots(x-(\alpha-1)),\quad{x\choose\alpha}=\frac{[x]_{\alpha}}{\alpha!},$
where $\alpha\geq 0$ is an integer and $x$ is a variable.
The sum for $\Phi_{i,r}$ can be expressed as
$\Phi_{i,r}=\sum_{\alpha_{1}+\cdots+\alpha_{n}=r}{r\choose\alpha_{1},\ldots,\alpha_{n}}a_{i,1}^{\alpha_{1}}\left(\frac{\partial}{\partial a_{i,1}}\right)^{\alpha_{1}}\cdots a_{i,n}^{\alpha_{n}}\left(\frac{\partial}{\partial a_{i,n}}\right)^{\alpha_{n}}.$
Here the integers $\alpha_{j}$ are nonnegative, and the multinomial appears in the sum because we need to choose $\alpha_{1}$ indices $j$ equal to $1$, $\alpha_{2}$ indices $j$ equal to $2$, and so on. Applying the one variable formula above, this simplifies to
$\Phi_{i,r}$ $=\sum_{\alpha_{1}+\cdots+\alpha_{n}=r}{r\choose\alpha_{1},\ldots,\alpha_{n}}[\partial_{1}]_{\alpha_{1}}\cdots[\partial_{n}]_{\alpha_{n}}$
$=\sum_{\alpha_{1}+\cdots+\alpha_{n}=r}r!{\partial_{1}\choose\alpha_{1}}\cdots{\partial_{n}\choose\alpha_{n}}.$
The claim is that this sum is equal to
$[\sum_{j}\partial_{j}]_{r},$
or after dividing by $r!$,
(${\sum_{j}\partial_{j}\choose r}=\sum_{\alpha_{1}+\cdots+\alpha_{n}=r}{\partial_{1}\choose\alpha_{1}}\cdots{\partial_{n}\choose\alpha_{n}}.$
This last equality clearly holds if we substitute any nonnegative integers into $\partial_{j}$.
Because the $\partial_{j}$ commute, both sides of the equality in the statement of the lemma can be expressed as polynomials in $\partial_{j}$ with integer coefficients. Since these polynomials take the same value when evaluated at the Zariski dense set of all nonnegative integers, they must be equal. Hence they also reduce to equal polynomials over any field. ∎
Proof of Theorem 2.11.
As in the proof of Theorem 1.2, we reduce to the case of the boundary of a simplex. We can write the degree map on $\Delta$ as a sum of the degree maps on boundaries of $d$-simplices $\Pi_{i}$, see (1). Let $\tilde{\Psi}_{r}$ be the differential operator on $K[x_{0},\dotsc,x_{n}]$ given by $\sum_{I}\frac{a^{I}}{I!}\partial^{I}$, where the sum is over all $d\times(n+1)$ matrices $I$ with all rows sum equal to $r$, i.e., $\tilde{\Psi}_{r}$ is given by the same formula as $\Psi_{r}$ except we include $\frac{\partial}{\partial{a_{0,j}}}$ as well. For each $i$, $\tilde{\Psi}_{r}\deg_{\Pi_{i}}(x^{J})$ is equal to the result of applying the differential operator which is defined only using the vertices of $\Pi_{i}$. This holds for $\Delta$ as well. We see that
$\Psi_{r}\deg_{\Delta}(x^{J})=\tilde{\Psi}_{r}\deg_{\Delta}(x^{J})=\tilde{\Psi}_{r}\deg_{\Pi_{1}}(x^{J})+\dotsb.$
If we know that $\tilde{\Psi}_{r}\deg_{\Pi_{i}}(x^{J})=(-1)^{dr}(r!)^{2d}\deg_{\Pi_{i}}(x^{J})$, then we deduce $\Psi_{r}\deg_{\Delta}(x^{J})=(-1)^{dr}(r!)^{2d}\deg_{\Delta}(x^{J})$ as well. We can therefore reduce to showing that $\Psi_{r}\deg(x^{J})=(-1)^{dr}(r!)^{2d}\deg(x^{J})$ when $\Delta=\Pi$ is the boundary of a simplex with vertex set $\{0,1,\dotsc,d\}$. We claim that, for each $1\leq i\leq d$, we have
$\phi_{i}\deg(x^{J})=-\deg(x^{J}).$
Indeed, cofactor expansion along the $i$th row implies that $\phi_{i}X_{j}=X_{j}$ for any $j$. The product rule then implies that $\phi_{i}$ acts on any Laurent monomial of degree $m$ in the $X_{j}$ by multiplication by $m$. By (2), $\deg(x^{J})$ is a Laurent monomial of degree $-1$ in the $X_{j}$. By Lemma 2.12, we have
$\begin{split}\Phi_{i,r}\deg(x^{J})&=\phi_{i}(\phi_{i}-1)\dotsb(\phi_{i}-(r-1))\deg(x^{J})\\ &=(-1)(-2)\dotsb(-r)\deg(x^{J})\\ &=(-1)^{r}r!\deg(x^{J}).\end{split}$
Applying (7) then gives the result. ∎